OpenStudy (anonymous):

Can some one help me solve this using calculator please !!! On a multiple choice test of 10 questions, each question has 5 possible answers.If 70% is the pass mark and Raj, who knows absolutely nothing about the subject, guesses each answer at random, determine the probability the he will pass ?

OpenStudy (anonymous):

yikes your math teacher must hate you this is going to take a while

OpenStudy (anonymous):

we can easily compute the probability that he gets all right that would be \[P(x=10)=\left(\frac{1}{5}\right)^{10}\]

OpenStudy (anonymous):

the probability he gets 9 right is \[P(x=9)=\binom{10}{9}\left(\frac{1}{5}\right)^9\left(\frac{4}{5}\right)^1\] does this ring a bell?

OpenStudy (anonymous):

the probability he gets 8 right is \[P(x=8)=\binom{10}{8}\left(\frac{1}{5}\right)^8\left(\frac{4}{5}\right)^2\] well only one more to go

OpenStudy (anonymous):

\[P(x=7)=\binom{10}{7}\left(\frac{1}{5}\right)^7\left(\frac{4}{5}\right)^3\]

OpenStudy (anonymous):

compute this suckers and add them up i would definitely use a wolfram since i haven't used a calculator for anything other than to compute averages in at least 4 yeras